You have a portfolio that is invested 15 percent in Stock R, 34 percent in Stock S, and the remainder in Stock T. The beta of Stock R is .60, and the beta of Stock S is 1.15. The beta of your portfolio is 1.26. What is the beta of the Stock T?
Answer #1

Solution:

Portfolio beta = (weight of R × Beta of R) + (weight of S × Beta of S) + (weight of T × Beta of T)

Weight of T = 100 - (weight of R + weight of S) = 100% - (15% + 34%) = 100% - 49% = 51%

1.26 = (15% × 0.60) + (34% × 1.15) + (51% × Beta of T)

1.26 = 0.09 + 0.391 + (51% × Beta of T)

1.26 - 0.481 = 51% × Beta of T

0.779 = 51% × Beta of T

Beta of T = 0.779 / 51%

Beta of T = 1.53.